car A and car B travel from town to a vacation cabin. car A travels at 45 mph. car B leaves two hours after car A and travels 15 mph faster than car A. if the cars arrive at the same time, how many miles is it from town to the cabin
Answer by mananth(16949)

time car A - time car B = 2 hours
Let distance to cabin be x
car A speed = 45
Car B speed = 45+15=60 mph
x/45 -x/60=2
LCD = 180
4x-3x= 2*180
x=360
x=360 miles